OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

use the formula of slope m=y2-y1/x2-x1 =0-4/2-(-6) m=-4/8 m=-1/2 for slop intercept take either point will give the same answer im using (-6,4) here x=-6 and y=4 put this in formula (4)=-1/2(-6)+b b=3-4 b=1 so y=-1/2x+1